Circular No. 5760

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 A. V. Filippenko and T. Matheson, University of California, Berkeley,

report: “We obtained spectra (range 300-1000 nm, resolution 0.5 nm) of SN 1993J on Apr. 13 UT with the Shane 3-m reflector at Lick Observatory. Well-developed P-Cyg profiles are visible in the hydrogen Balmer lines, Fe II, Na D, Mg II, and Ca II. A component of moderate width (FWHM = 3.4 nm) is superposed on the broad Hα emission (FWHM = 29 nm). The narrow emission lines of He II 468.6 nm, [Fe X] 637.4 nm, and Hα previously noted on IAUC 5740 are now gone; the ionized circumstellar gas that produced them has now presumably been overcome by the supernova ejecta. This may be consistent with the appearance of radio and x-ray emission. The narrow absorption line at 373.45 nm, tentatively identified as Al II on IAUC 5740, is not visible; either it was an artifact, or it was produced by the circumstellar gas.”

S. Odewahn, University of Minnesota; C. E. Woodward, University of Wyoming; and J. Bailey, Agnes Scott College, report the following infrared magnitudes of SN 1993J, obtained with the 1.3-m KPNO telescope (+ SQIID imager): Apr. 1.21 UT, J = 10.54 ± 0.02, H = 10.47 ± 0.02, K = 10.37 ± 0.02; Apr. 2.31, 10.62 ± 0.01, 10.49 ± 0.01, 10.33 ± 0.01. Photometric magnitudes for the reference stars in the supernova field are as follows: GSC 4383.0565, J = 10.36 ± 0.01, H = 10.14 ± 0.01, K = 10.09 ± 0.01; GSC 4383.0928, 10.85 ± 0.01, 10.63 ± 0.01, 10.57 ± 0.02; GSC 4383.0340, 13.32 ± 0.05, 13.12 ± 0.05, 13.03 ± 0.05.

Further photoelectric photometry (cf. IAUC 5750, 5755, 5758) by D. Hanzl, N. Copernicus Planetarium, Brno (H); by T. Kato, Kyoto University (K); by M. Okyudo and T. Ishida, Nishi-Harima Observatory (OI); by M. Mikołajewski and K. Rumiński, N. Copernicus University, Toruń (MR); by R. Mutel and P. Sauerbrei, University of Iowa (MS); and by P. Prugniel, Observatoire de Haute Provence (P): Apr. 8.576 UT, V = 11.68 (OI); 9.541, V = 11.46 (K); 9.659, V = 11.52 (OI); 10.14, B = 11.81 ± 0.06, V = 11.42 ± 0.06, R = 10.81 ± 0.05, I = 10.70 ± 0.10 (MS); 10.588, V = 11.46 (OI); 10.91, V = 11.08 ± 0.08, B–V = +0.53 ± 0.10 (H); 10.92, V = 11.26 ± 0.01, U–B = –0.17 ± 0.03, B–V = +0.46 ± 0.01, V–R = +0.51 ± 0.01, V–I = +0.36 ± 0.02 (MR); 12.00, B_{c} = 11.85, B_{c}–R_{c} = +0.83 (P); 12.476, V = 11.11 (K); 12.499, I_{c} = 10.71 (K); 12.537, B = 11.57 (K).

1993 April 13 Daniel W. E. Green